OpenStudy (anonymous):

find the general solution of the differential equation y'=4+2x+2y+xy

OpenStudy (anonymous):

y'-y(2+x)=2x+4

OpenStudy (anonymous):

check my algebra

OpenStudy (anonymous):

from here it is just First Order Linear Differential Equation

OpenStudy (anonymous):

i don't see how u got it to that form

OpenStudy (anonymous):

yes i do sorry

OpenStudy (anonymous):

find integrating factor \[e^{\int 2+x}\]

OpenStudy (aravindg):

dashini help

OpenStudy (anonymous):

yea i'm there and just got that

OpenStudy (anonymous):

\[y(x) = c_1 e ^{((x^2/2)+2 x)}-2\]

OpenStudy (anonymous):

This is your integrating factor \[e^{\left(2x+\frac{x^2}{2}\right)}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now put that back into the original equation and solve it for y correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay great...thank you

OpenStudy (anonymous):

\[\left(y e^{\left(2x+\frac{x^2}{2}\right)}\right)'=(2x+4)\left( e^{\left(2x+\frac{x^2}{2}\right)}\right)\]

OpenStudy (anonymous):

\[\left(y e^{\left(2x+\frac{x^2}{2}\right)}\right)=\int (2x+4)\left( e^{\left(2x+\frac{x^2}{2}\right)}\right) \text{dx}\]

OpenStudy (anonymous):

\[\left(y e^{\left(2x+\frac{x^2}{2}\right)}\right) =2 e^{\frac{1}{2} x (4+x)}+C\]

OpenStudy (phi):

Though it is not obvious you can factor 4 + 2x +2y +xy into x(y+2) +2(y+2) = (x+2)(y+2) and now you have a separable equation dy/(y+2) = (x+2)dx

OpenStudy (anonymous):

i like that way better :)

OpenStudy (nikvist):

\[y'=4+2x+2y+xy=(x+2)(y+2)\]\[\frac{y'}{y+2}=(\ln(y+2))'=x+2\]\[\ln{(y+2)}=\frac{x^2}{2}+2x+C\]\[y=e^{x^2/2+2x+C}-2=C_1e^{x^2/2+2x}-2\]

OpenStudy (phi):

The first way is more general.

OpenStudy (anonymous):

yeah, much easier using separable equation

OpenStudy (anonymous):

now with the condition of y(0) = 0 does this change everything?

OpenStudy (anonymous):

nope, you just get value for C